# Evidence for Low Universal Equilibrium Black Hole Spin in Luminous Magnetically Arrested Disks

### v1

@@ 3.3.1 Weaker spin-down in thin disks
is the absolute BH magnetic flux, and the integral is taken over the BH event horizon. We observe that thicker disks have a larger normalized magnetic flux at the event horizon than thinner disks, consistent with previous findings (Avara et al., 2016; Scepi et al., 2023; Dhang et al., 2024). Additionally, we note that $\phi_{\rm BH}$ in thin and thick disks exhibits distinct spin dependencies. Whereas thin disks show a somewhat monotonic trend with spin, thick disks display a clear maximum around $a\sim 0.3$.
We now examine whether the weaker magnetic flux in thin MADs is directly reflected in their jet power. Via the Blandford-Znajek (BZ, Blandford & Znajek, 1977) mechanism, the EM jets are powered by the BH rotational energy, which leads to the following scaling for the jet power (Tchekhovskoy et al., 2010),
$P_{\rm BZ}=\frac{\kappa}{4\pi c}\Omega^{2}_{\rm H}\Phi^{2}_{\rm BH}f(\Omega_{\rm H}),$ (23)
where $\kappa$ is a constant that depends on the magnetic field geometry, $c$ is the speed of light, $\Omega_{\rm H}=a/r_{\rm H}$ is the BH rotational frequency. Here $\Phi_{\rm BH}$ is the magnetic flux threading the BH event horizon, and $f(\Omega_{\rm H})=1+1.38{(\Omega_{\rm H}r_{\rm g}/c)}^{2}-9.2{(\Omega_{\rm H}r_{\rm g}/c)}^{4}$ is a high-spin correction. Here we use the value $\kappa=0.054$ for a jet with a split-monopole geometry.
We observe in Eq. (23) that the BZ jet power is a function of $\Phi^{2}_{\rm BH}$, $\kappa$ and $a$. To explain why jets are less powerful in luminous MADs than in nonradiative MADs, the bottom panel of Figure 6 shows the EM power normalized by $\frac{1}{4\pi}\kappa\Phi_{\rm BH}^{2}$, with $\kappa=0.054$, isolating the spin dependence of the jet power. Here, we measure the EM (jet) power $P_{\rm EM}$ as the EM component of the total energy flux, $\dot{E}(r)$ (Eq. 16, on the BH (not normalized to the mass flux).
Surprisingly, once normalized to the magnetic flux on the BH, the EM power becomes the same for all disk thermal scale heights. Thus, it is $\Phi_{\rm BH}$ that sets the jet power, with thinner disk having lower values of $\Phi_{\rm BH}$. However, what sets $\Phi_{\rm BH}$ or $\phi_{\rm BH}$ is largely unknown.
Now we look at the angular momentum extraction due to the EM jets. As we saw in Figure 4, $l_{\rm EM}$ has a steep dependence on the thermal scale height, changing substantially when going from $h/r=0.3$ to $h/r=0.1$. Following Moderski & Sikora (1996) and Lowell et al. (2024), we can describe the EM specific angular momentum flux on the BH as $l_{\rm EM}=\eta_{\rm EM}/(k\Omega_{\rm H})$ to better understand its behavior.
In Figure 7 we show the ”slippage” parameter, $k$, as a function of spin for the different disk thermal thicknesses. We see that the thinner the disk, the larger the value of $k$. A higher $k$, indicating faster rotation of field lines, results in less efficient angular momentum extraction and thus a lower $l_{\rm EM}$. For a MAD of a given $h/r$, this leads to a higher equilibrium spin.
Figure 7: $k=\Omega_{\rm F}/\Omega_{\rm H}$ as a function of BH spin. The gray points show the values calculated in Lowell et al. (2024), and the colored points show cooled disks of various scale heights. For any value of spin, thick, radiatively-inefficient disks have jets with the lowest angular frequency. As the disk cools, the value of $k$ increases. $k$ appears to converge as the disk becomes thinner.
Figure 8: Last magnetic field line attached to BH with spin $a=0.9375$ for three simulations of varying scale height. We normalize the radial coordinate to the Alfvén radius, $R_{\rm A}=r_{g}/\Omega_{\rm F}$, also shown by the vertical blue line. For a distance greater than roughly $R_{\rm A}\sim 8r_{\rm g}$ we find that the thinner the disk, the wider the jet opening angle. See Figure 7, where thin disks have a larger value of $k=\Omega_{\rm F}/\Omega_{\rm H}$, implying that jets with a wider opening angle have jets that are rotating faster than more collimated jets. We notice that both thick disks are parallel to each other and therefore have similar $\nu$ while the thick disk has a larger $\nu$. We show the high resolution simulation for $h/r=0.3$ (H3a0.94hr).
@@
The rotation velocity of the magnetic field lines, $k$, is tied to the jet’s structure, making it useful to examine the magnetic field lines connected to the BH in order to understand BH spin-down. In Figure 8 we plot the outermost magnetic field line attached to the BH averaged in time and $\varphi$. This field line represents the jet shape at the boundary where it interacts with the wind. One can approximate the flux of a collimating field line as a power-law in distance (Tchekhovskoy et al., 2010),
@@
where $\nu$ sets the jet shape: $\nu=0$ corresponds to monopolar and $\nu=1$ to parabolic jet. In a realistic jet, the flux will not follow such a simple power-law. Nevertheless, we can use this approximation to understand the relationship between the jet shape and jet angular momentum flux. From Eq. (24), it follows that
$\sin\left(\frac{\theta}{2}\right)=\sqrt{\frac{a_{\varphi}}{2}}\left(\frac{r}{R_{\rm A}}\right)^{-\nu/2},$ (25)
@@
where $\theta$ is the polar angle, $a_{\varphi}$ is the enclosed magnetic flux, constant along the field line, $r$ is the distance from the black hole, and $\nu$ is a parameter describing the jet shape that we generalize to depend on radius. We normalize the distance to the Alfvén radius, $R_{\rm A}=1/\Omega_{\rm F}$, where the Alfvén velocity is equal to the outward velocity of the gas. At the Alfvén surface the outflowing material can no longer communicate back to the even horizon through Alfvén waves. Indeed, only the jet shape inside of the Alfvén surface can affect the jet launching and hence, the value of $k$. Therefore, the Alfvén surface defines a meaningful radius where we can compare the value of $\nu$ for different simulations.
@@
We show jet shapes for models with $a=0.9375$ of $h/r=0.3,0.1,0.03$ using gray, purple, and orange curves, respectively. For each model, we use the value of $k$ (shown in Figure 7) to compute $R_{\rm A}$.
@@
We find that the thinner the accretion disk, the wider the jet barrel at the Alfvén radius. The value of $\nu$ at $R_{\rm A}$ for the thick disk is close to $\sim 1$, whereas the value for the two thinner disks ($h/r=0.1$ and $h/r=0.03$) is around $\sim 1/4$. We can also see this from Figure 1: disks with smaller thermal thickness collimate jets less than thicker disks. The thinner the disk, the greater the jet opening angle, with magnetic field lines increasingly resembling a monopole geometry ($\nu=0$) as scale height decreases. Thinner disks allow more space for jet widening, and MHD acceleration becomes most efficient with larger lever arms. As a result, field lines rotate more rapidly (with a larger $\Omega_{\rm F}$ and consequently greater $k$, as seen in Figure 7).
@@
This picture is consistent with the work of Tchekhovskoy et al. (2010) who studied $k$ as a function of the index $\nu$, and found that different jet geometries lead to different values of $k$. A monopole ($\nu=0$) leads to a roughly constant $k\sim 0.5$, while a parabolic geometry ($\nu=1$) leads to a $k$ that decreases with angle and has an average value [^2] We note that, in Tchekhovskoy et al. (2010) $k$ never dropped below $0.25$. of $\sim 0.4$. We conclude that larger time-averaged values of $k$ are consistent with jet shapes approaching a monopole as $h/r$ decreases. We also stress that the change in $k$ and jet power is the strongest between $h/r=0.3$ and $h/r=0.1$. The change between $h/r=0.1$ and $h/r=0.03$ is almost imperceptible, which is consistent with their slopes being the same at the Alfvén surface.
3.3.2 Decoupling of thermal and magnetic dynamics
As the thermal scale height decreases in Figure 3, $a_{\rm eq}$ appears to converge. This could imply that for even thinner disks, the equilibrium spin might continue to asymptotically approach a relatively small value of spin $a\lesssim 0.31$, much lower than the analytic value for NT razor-thin disks, $a=0.998$. For the thermal scale height to drop from $h/r=0.1$ to $h/r=0.03$ and leave the spin evolution unchanged, some physical property of the disk structure needs to remain independent of the scale height at $h/r\lesssim 0.1$. As we show below, the magnetic structure and magnetic forces are roughly independent of the thermal scale height for $h/r\lesssim 0.1$.
In thin MADs, magnetic pressure—both its large-scale and turbulent components—controls the vertical equilibrium of the accretion disk, and thermal pressure becomes highly sub-dominant. This may explain why BH spin-down by thin MADs is unaffected by the thermal $h/r$ value: this is because the primary torque on the BH comes from the jet’s magnetic fields, which appear to remain unchanged for $h/r\leq 0.1$.
@@ 3.3.2 Decoupling of thermal and magnetic dynamics
We begin by examining the thermal and magnetic forces governing disk compression and expansion in the latitudinal ($\theta$) direction, disregarding centrifugal force since it is unrelated to jet formation, BH energy and angular momentum extraction, or thermal $h/r$. While thermal and radiation pressures act to expand the disk, magnetic pressure plays a more complex role. Turbulent and large-scale (laminar) magnetic pressures contribute differently to the disk dynamics, influencing its overall equilibrium in distinct ways. Turbulent magnetic pressure arises from disk turbulence, providing support against collapse, whereas large-scale (laminar) magnetic pressure results from fields anchored to the jet structure, compressing the disk. One can separate both terms using a Reynolds decomposition,
@@
We show the $\theta$-profiles of the magnetic pressures and the thermal pressure in Figure 9 for the model H1a0.94. We see that the thermal and turbulent magnetic pressures have similar vertical profiles, whereas the large-scale laminar magnetic pressure has an inverted profile. The reason for this difference is that $\langle\delta b^{2}\rangle_{t,\varphi}$ and $\langle b\rangle^{2}_{t,\varphi}$ play different roles  [^3] Explicit compression and expansion can be understood by examining the inverse of the gradient of the various terms in Fig. 9. A positive gradient, $P(\theta_{2})-P(\theta_{1})>0$ where $\theta_{1}=\pi/2$ and $|\theta_{2}-\pi/2|>0$, results in compression, while a negative gradient, $P(\theta_{2})-P(\theta_{1})<0$, leads to expansion.: (1) the laminar magnetic pressure from the jet compresses the disk and (2) the turbulent magnetic pressure, along with the thermal pressure, resists the disk compression and maintains the magnetohydrostatic equilibrium. In radiative simulations, the radiation pressure also resists disk compression, as expected. The compressive role of $\langle b\rangle^{2}_{t,\varphi}$ is an established result from self-similar theory (Ferreira & Pelletier, 1995), and the role of the turbulent magnetic pressure on the vertical equilibrium is consistent with previous 3D (GR)MHD simulations (Salvesen et al., 2016; Jacquemin-Ide et al., 2021; Scepi et al., 2023). Similar vertical pressure profiles are observed across all simulations at all radii, modulated by disk thickness; thicker disks exhibit broader pressure profiles. However, the relative magnitudes of the various pressures depend on disk thickness.
@@
We now focus on how the relative magnitudes of the pressures depend on the disk thickness. In Figure 10 we plot the ratio of the turbulent magnetic pressure to the sum of the thermal and radiation pressures in the disk mid-plane for three simulations with $a=0.9375$ and $h/r=0.3$, $0.1$, and $0.03$, shown as gray, purple, and orange lines, respectively. Since these pressures all act to expand the disk, this ratio indicates the dominant pressure support mechanism for the accretion disk. A clear dichotomy emerges between thick ($h/r=0.3$) and thin disks ($h/r=0.1$ and $h/r=0.03$). Thin disks are magnetically dominated at the midplane over a larger range of radii, $r\lesssim 15\,r_{\rm g}$ compared to thick disks, which are only magnetically dominated at $r\lesssim 3\,r_{\rm g}$. We conclude that thin disks are supported primarily by turbulent magnetic pressure, whereas thick disks are thermally supported, aligning with previous findings (Scepi et al., 2023). This shows that the disk thermal scale height is now decoupled from the disk vertical equilibrium and thus is not a relevant quantity for characterizing the system.
@@
Figure 9: Vertical profiles at $r=6r_{\rm g}$ of time-averaged magnetic pressures for H1a0.94 normalized to thermal pressure in the mid-plane. Large-scale laminar magnetic pressure (magenta) dominates over turbulent magnetic pressure (purple) and thermal pressure (orange) above and below the disk. Turbulent magnetic pressure dominates the disk midplane.
@@
Figure 10: Ratio of turbulent magnetic pressure to thermal and radiation pressures measured in the equatorial plane for varying disk scale heights with BH spin of $a=0.9375$. The thick MAD (gray) is dominated by magnetic pressure in the inner few radii, but thermal pressure dominates beyond $4r_{g}$. The cooled MAD (purple) is dominated by magnetic pressure overall. Magnetic pressure also dominates in Ra0.94, increasing with radius until $\frac{1}{2}\langle\delta b^{2}\rangle/(p_{\rm g}+p_{\rm rad})>6$ at $10r_{g}$. We show the high resolution simulation for $h/r=0.3$ (H3a0.94hr).
@@
In Section 3.3.1, we demonstrated that jet shape (parabolic or monopolar) directly determines the $k$ factor, playing a key role in regulating the jet angular momentum flux from the BH. We also found that simulations with $h/r=0.1$ and $h/r=0.03$ exhibit identical jet shape slopes, both forming monopolar-like jets, whereas thick disks produce parabolic jets. Geometric arguments suggest a connection between jet morphology and disk thermal thickness, which we propose is mediated by the compression of the disk by the large-scale magnetic field attached to the jet.
@@
$F_{\rm b,{jet}}=-\frac{1}{\langle\rho\rangle_{t,\varphi}}\partial_{\theta}\left(\sqrt{-g}\langle b\rangle^{2}_{t,\varphi}/2\right).$ (27)
@@
We divide by the local density, as magnetic acceleration and compression strength depend on density, with lower densities enabling more efficient acceleration. Curvature terms are ignored, as they are found to be negligible. We also ignore the vertical ($\theta$) magnetic tension forces, as we also find them to be negligible (this is not the case for radial tension forces, see Chatterjee & Narayan, 2022).
@@
In Figure 11, we present $F_{\rm b,{jet}}$, the density-normalized compression force measured at $r=3\,\,r_{g}$, for four simulations with $a=0.9375$: $h/r=0.3$, $h/r=0.1$, $h/r=0.03$ 2T, and $h/r=0.03$ 1T shown as gray, purple, orange and black lines, respectively. Arrows indicate the force direction: a positive force pushes along $\theta$, while a negative force pushes in the opposite direction. All simulations exhibit similar profiles, showing compression towards the disk midplane above and below it and an expansion region above and below the compression zone. We associate the expansion region with jet expansion and confinement; however, this is not our region of interest. The magnitude of disk compression reveals a dichotomy between the thin ($h/r=0.1$ and $h/r=0.03$) and thick ($h/r=0.3$) disks: thicker disks display stronger compression above and below the disk midplane. Furthermore, all simulations of thin disks show roughly similar disk compression magnitudes, even though their disk thicknesses differ by a factor of 3.
@@
We find stronger disk compression for thicker disks, because thicker disks restrict the jet expansion, resulting in a stronger backreaction against the disk. When the disk becomes sufficiently thin ($h/r\leq 0.1$), further thinning has minimal impact since the jet occupies most of the available space. The twofold increases in $F_{\rm b,{jet}}$ compression for thicker vs. thinner disks is consistent with the twofold increase in jet efficiency observed in Fig. 6 for thicker vs. thinner disks  [^4] The fact that $F_{\rm b,{jet}}$ is strongest for thick disks may seem at odds with thinner disks being more highly magnetized at the midplane (see Fig. 10). However, the ratio $\frac{1}{2}\langle\delta b^{2}\rangle/(p_{\rm g}+p_{\rm rad})$ can be misleading for comparing magnetic field strength across simulations, as thermal (and radiative) pressure also changes with $h/r$. .

### v2

@@ IV Semi-analytic model for BH spin-down in thin MADs
Figure 6(a) shows that our above model for $\eta_{\mathrm{EM}}^{\mathrm{thin}}$ well reproduces the jet efficiency for thin MADs.
We fit the spin dependence of the $k$ factor using the data shown in Fig. 7, with a power law, [^3] In this definition, we set $k(a=0)=0.2$, though strictly speaking, $k(a=0)$ is undefined. This choice, made for the quality of the fit, does not affect our results since the angular momentum extraction term goes to zero when $a=0$, as $\eta_{\mathrm{EM}}(a=0)=0$.
[…]
Figure 4, shows that our fit for $l_{\mathrm{EM}}=\eta_{\mathrm{EM}}^{\mathrm{thin}}/{k_{\mathrm{thin}}\Omega_{\mathrm{H}}}$, as a function of $a$, reproduces the simulation data well. By combining the above EM and HD components of $l$ and $e$, and replacing $\eta_{\mathrm{EM}}^{\mathrm{thin}}$ with $\eta_{\mathrm{EM}}^{\mathrm{thick}}/2$ in Eq. (23), we obtain the full thin MAD semi-analytic model,
$\begin{split}s_{\mathrm{MAD}}^{\mathrm{thin}}=0.4l_{\mathrm{NT}}-0.85e_{\mathrm{NT}}\times 2a-\frac{\eta_{\mathrm{EM}}^{\mathrm{thick}}}{2}\left(\frac{1}{k_{\mathrm{thin}}\Omega_{\mathrm{H}}}-2a\right),\end{split}$ (32)
for BH spin-down in thin MADs. Figure 2 shows our spin evolution model (Eq. 32) with the purple dashed line, which reproduces the simulation data well. Although our model is tailored to $h/r=0.1$ MADs, it also well reproduces the behavior of $s$ for thinner MADs, thanks to the universality of thin MAD structure (see Section V.2).
Figure 8 shows BH spin evolution as a function of mass accreted onto the BH for thin MADs with $h/r=0.1$. For this, we solve the coupled ODEs, Equations (2) and (4), and use spline interpolations for the specific energy and angular momentum fluxes at the horizon, for simulations (solid lines) with $h/r=0.1$ (models H1a#). We show solutions for the initial spin values $a_{0}=-1$, 0, $0.5$, and $1$. We find that regardless of the initial spin, $a$ evolves to an equilibrium value of $a_{\mathrm{eq}}\approx 0.3$. Most initial spins get within 10% of the equilibrium value after less than 85% of the initial BH mass is accreted, while $a_{0}=-1$ requires slightly more, roughly 100% of the initial mass. Thin MADs with $a=1$ spin down to $a=0.5$ after accreting 25% of the initial BH mass, a slower process compared to thick MADs, which only require the accretion of 10% of the BH mass for the same spin-down. We also solve the ODEs using our thin MAD model (Equation 32, dashed lines) and confirm that our model reproduces the simulation (solid lines) well. We associate the reduced BH spin-down efficiency in thin MADs with a general decrease in angular momentum and energy extraction by the jets (see Section III.3).
Figure 8: Our semi-analytic model (dashed lines, Eq. 32) accurately describes BH spin evolution (solid lines) in thin MADs. We show the BH spin evolution, $a(m/M_{0})$, for thin MADs with $h/r=0.1$ (models H1a#) as a function of $m/M_{0}$, the ratio of accreted mass to the initial mass of the BH. When $m/M_{0}=1$, the BH has accreted its own (initial) mass. The upper $x$-axis gives spin evolution in the units of Myr over an initial Eddington rate, $\lambda_{0}$. We show solutions for the initial spin values, $a_{0}=-1$, 0, $0.5$, and $1$: for any initial spin, the BH evolves to an equilibrium spin, $a_{\mathrm{eq}}\approx 0.29$. For an initial BH spin of $a=1$, a thin MAD will spin down its BH to $a=0.5$ after accreting $25\%$ of its initial mass. For most initial spins, the BH must accrete around $85\%$ to reach equilibrium spin. For negative initial spin, the BH must accrete more mass than for the positive initial spins to reach $a_{\mathrm{eq}}$, roughly $125\%$ of $M_{0}$. Dashed lines, which show the spin evolution using our semi-analytic model (Eq. 32), agree well with the thin MAD spin-down model, which uses spline fits to $e_{\mathrm{in}}$ and $l_{\mathrm{in}}$.
V Jet magnetic structure and forces
As the thermal scale height decreases in Figure 3, $a_{\mathrm{eq}}$ appears to converge to a universal value. This could imply that for even thinner disks, the equilibrium spin might continue to asymptotically approach a relatively small value of spin $a\lesssim 0.31$, much lower than the analytic value for NT razor-thin disks, $a=0.998$. For the thermal scale height to drop from $h/r=0.1$ to $h/r=0.03$ and leave the spin evolution unchanged, some physical property of the disk structure needs to remain independent of the scale height at $h/r\lesssim 0.1$.
In this section, we show that the magnetic structure–including the jet’s magnetic structure and the magnetic forces that squeeze the accretion disk–remains independent of scale height for $h/r\lesssim 0.1$. First, in section V.1, we demonstrate that thin disks ($h/r\lesssim 0.1$) have monopolar jet morphologies, while thick disks ($h=0.3$) exhibit parabolic morphologies. Second in section V.2, we show that for $h/r\lesssim 0.1$, the forces influencing disk equilibrium become predominantly magnetic. As a result, the disk structure becomes independent of the thermal scale height, leading to little or no change in the magnetic forces between models with $h/r=0.1$ and $h/r=0.03$.
This may explain why BH spin-down by thin MADs is unaffected by the thermal $h/r$ value: this is because the primary torque on the BH comes from the jet’s magnetic fields, which appear to remain unchanged for $h/r\leq 0.1$.
V.1 Jet’s structure
@@
In this section we look at the jet’s structure to see how it changes with $h/r$. In Figure 9 we plot the time- and $\varphi$-average shape of the outermost magnetic field line attached to the BH. This field line represents the jet shape at the boundary where it interacts with the wind. One can approximate the magnetic flux of a collimating field line as a power law in distance [94],
@@
where $\theta$ is the polar angle, $a_{\varphi}$ is the enclosed magnetic flux divided by $2\pi$ and constant along the field line, $r$ is the spherical polar distance from the center of the black hole, and $\nu$ is a parameter describing the jet shape that we generalize to depend on radius. The parameter $\nu$ sets the jet shape: $\nu=0$ corresponds to monopolar and $\nu=1$ to parabolic jets. In a realistic jet, the flux will not follow such a simple power-law. Nevertheless, we can use this approximation to understand the relationship between the jet shape and jet angular momentum flux.
Figure 9: Thick MADs produce more collimated jets than thin MADs. Top panel: the shape of the outermost magnetic field line attached to BH with spin $a=0.9375$ for three simulations of varying scale height (here, we use the data for the $h/r=0.3$ HRHR simulation, H3a0.94hr). We normalize the radial coordinate to the Alfvén radius, $R_{\mathrm{A}}=r_{\mathrm{g}}/\Omega_{\mathrm{F}}$, also shown by the vertical dashed blue line. For a distance greater than roughly $R_{\mathrm{A}}\sim 8r_{\mathrm{g}}$ we find that the thinner the disk, the wider the jet opening angle. (See also Figure 7, where thin disks have a larger value of $k=\Omega_{\mathrm{F}}/\Omega_{\mathrm{H}}$, implying that less collimated jets rotate faster than more collimated ones.) Bottom panel: The $\nu$ jet shape parameter of the field lines in the top panel. The jets in the thick MAD (gray) more closely follows a parabolic geometry, while jets in thin MADs (purple for $h/r=0.1$ and orange for $h/r=0.03$) are closer to a monopole geometry. This is consistent with the top panel, in which both thin disk dependencies are parallel to each other and therefore have similar $\nu$ values, whereas the thick disk has a larger $\nu$ value.
From Eq. (33), it follows that
$\sin\left(\frac{\theta}{2}\right)\propto\left(\frac{r}{R_{\mathrm{A}}}\right)^{-\nu/2},$ (34)
@@
We normalize the distance to the Alfvén radius, $R_{\mathrm{A}}=1/\Omega_{\mathrm{F}}$, where the Alfvén velocity is equal to the outward velocity of the gas. At the Alfvén surface the outflowing material can no longer communicate back to the event horizon through Alfvén waves. Indeed, only the jet shape inside of the Alfvén surface can affect the jet launching and hence, the value of $k$. Therefore, the Alfvén surface defines a meaningful radius where we can compare the values of $\nu$ for different simulations.
@@
We show jet shapes for models with $a=0.9375$ of $h/r=0.3,0.1,0.03$ using gray, purple, and orange curves, respectively. For each model, we use the value of $k$ (shown in Figure 7) to compute $R_{\mathrm{A}}$.
@@
We find that the thinner the accretion disk, the wider the jet barrel at the Alfvén radius. We can also see this from Figure 1: disks with smaller thermal thickness collimate jets less than thicker disks. The thinner the disk, the greater the jet opening angle, with magnetic field lines increasingly resembling a monopole geometry ($\nu=0$) as the disk scale height decreases. Thinner disks allow more space for jet widening, and MHD acceleration becomes most efficient with larger lever arms. As a result, field lines rotate more rapidly (with a larger $\Omega_{\mathrm{F}}$ and consequently greater $k$, as seen in Figure 7).
In the bottom panel of Figure 9 we show the $\nu$ parameter, obtained by solving for the slope in the top panel. The value of $\nu$ at $R_{\mathrm{A}}$ for the thick disk is $\sim 0.6$, whereas the values for the two thinner disks ($h/r=0.1$ and $h/r=0.03$) are distinctly lower, both around $\sim 0.38$.
@@
This picture is consistent with the work of Tchekhovskoy et al. [94] who studied $k$ as a function of the index $\nu$, and found that different jet geometries lead to different values of $k$. A monopole ($\nu=0$) leads to a roughly constant $k\sim 0.5$, while a parabolic geometry ($\nu=1$) leads to a $k$ that decreases with angle and has an average value [^4] We note that, in Tchekhovskoy et al. [94] $k$ never dropped below $0.25$. of $\sim 0.4$. We conclude that larger time-averaged values of $k$ are consistent with jet shapes approaching a monopole as $h/r$ decreases. We also stress that the change in $k$ and jet power is the strongest between $h/r=0.3$ and $h/r=0.1$. The change between $h/r=0.1$ and $h/r=0.03$ is almost imperceptible, which is consistent with their slopes being the same at the Alfvén surface.
V.2 Decoupling of thermal and magnetic forces
@@
In this section we study the forces governing the disk equilibrium and how they change with $h/r$. We examine the thermal and magnetic forces governing disk compression and expansion in the latitudinal ($\theta$) direction, disregarding centrifugal force since it is unrelated to jet formation, BH energy and angular momentum extraction, or thermal $h/r$. While thermal and radiation pressures act to expand the disk, magnetic pressure plays a more complex role. Turbulent and large-scale (laminar) magnetic pressures contribute differently to the disk dynamics, influencing its overall equilibrium in distinct ways. Turbulent magnetic pressure arises from disk turbulence, providing support against collapse, whereas large-scale (laminar) magnetic pressure results from fields anchored to the jet structure, compressing the disk. One can separate both terms using a Reynolds decomposition,
@@
We show the $\theta$-profiles of the magnetic pressures and the thermal pressure in Figure 10 for the model H1a0.94. We see that the thermal and turbulent magnetic pressures have similar vertical profiles, whereas the large-scale laminar magnetic pressure has an inverted profile. The reason for this difference is that $\langle\delta b^{2}\rangle_{t,\varphi}$ and $\langle b\rangle^{2}_{t,\varphi}$ play different roles  [^5] Explicit compression and expansion can be understood by examining the inverse of the gradient of the various terms in Fig. 10. A positive gradient, $P(\theta_{2})-P(\theta_{1})>0$ where $\theta_{1}=\pi/2$ and $|\theta_{2}-\pi/2|>0$, results in compression, while a negative gradient, $P(\theta_{2})-P(\theta_{1})<0$, leads to expansion.: (1) the laminar magnetic pressure from the jet compresses the disk and (2) the turbulent magnetic pressure, along with the thermal pressure, resists the disk compression and maintains the magnetohydrostatic equilibrium. In radiative simulations, the radiation pressure also resists disk compression, as expected. The compressive role of $\langle b\rangle^{2}_{t,\varphi}$ is an established result from self-similar theory [19], the role of the turbulent magnetic pressure on the vertical equilibrium is consistent with previous 3D (GR)MHD simulations [83, 34, 84], and more recent self-similar models [104]. Similar vertical pressure profiles are observed across all simulations at all radii, modulated by disk thickness; thicker disks exhibit broader pressure profiles. However, the relative magnitudes of the various pressures depend on disk thickness.
@@
We now focus on how the relative magnitudes of the pressures depend on the disk thickness. In Figure 11 we plot the ratio of the turbulent magnetic pressure to the sum of the thermal and radiation pressures in the disk midplane for three simulations with $a=0.9375$ and $h/r=0.3$, $0.1$, and $0.03$, shown as gray, purple, and orange lines, respectively. Since these pressures all act to expand the disk, this ratio indicates the dominant pressure support mechanism for the accretion disk. A clear dichotomy emerges between thick ($h/r=0.3$) and thin disks ($h/r=0.1$ and $h/r=0.03$). Thin disks are magnetically dominated at the midplane over a larger range of radii, $r\lesssim 15\,r_{\mathrm{g}}$ compared to thick disks, which are only magnetically dominated at $r\lesssim 3\,r_{\mathrm{g}}$. We conclude that thin disks are supported primarily by turbulent magnetic pressure, whereas thick disks are thermally supported, in agreement with previous findings [84]. This shows that the disk thermal scale height is now decoupled from the disk vertical equilibrium and thus is not a relevant quantity for characterizing the system.
@@
Figure 10: Vertical profiles at $r=6r_{\mathrm{g}}$ of time-averaged magnetic pressures for H1a0.94 normalized to thermal pressure in the midplane. Large-scale laminar magnetic pressure (magenta) dominates over turbulent magnetic pressure (purple) and thermal pressure (orange) above and below the disk. Turbulent magnetic pressure dominates the disk midplane.
@@
Figure 11: Ratio of turbulent magnetic pressure to thermal and radiation pressures measured in the equatorial plane for varying disk scale heights with BH spin of $a=0.9375$. The thick MAD (gray) is dominated by magnetic pressure in the inner few radii, but thermal pressure dominates beyond $4r_{\mathrm{g}}$. The thin MAD (purple) is dominated by magnetic pressure overall. Magnetic pressure also dominates in Ra0.94, increasing with radius until $\frac{1}{2}\langle\delta b^{2}\rangle/(p_{\mathrm{g}}+p_{\mathrm{rad}})>6$ at $10r_{\mathrm{g}}$. We show the high-resolution simulation for $h/r=0.3$ (H3a0.94hr).
@@
In Section V.1, we demonstrated that jet shape (parabolic or monopolar) directly determines the $k$ factor, playing a key role in regulating the jet angular momentum flux from the BH. We also found that simulations with $h/r=0.1$ and $h/r=0.03$ exhibit identical jet shape slopes, both forming monopolar-like jets, whereas thick disks produce parabolic jets. Geometric arguments suggest a connection between jet morphology and disk thermal thickness, which we propose is mediated by the compression of the disk by the large-scale magnetic field attached to the jet.
@@
$F_{\mathrm{b,{jet}}}=-\frac{1}{\sqrt{-g}\langle\rho\rangle_{t,\varphi}}\partial_{\theta}\left(\sqrt{-g}\langle b\rangle^{2}_{t,\varphi}/2\right).$ (36)
@@
We divide by the local density, as magnetic acceleration and compression force strength depend on density, with lower densities enabling more efficient acceleration. We ignore the curvature terms, because we found them to be negligible. We also ignore the vertical ($\theta$-) magnetic tension forces, as we also found them to be negligible [this is not the case for radial tension forces, see 6].
@@
In Figure 12, we present $F_{\mathrm{b,{jet}}}$, the density-normalized compression force measured at $r=3r_{\mathrm{g}}$, chosen as a representative radius close to the BH. We plot $F_{\mathrm{b,{jet}}}$ for four simulations with $a=0.9375$: $h/r=0.3$, $h/r=0.1$, $h/r=0.03$ 2T, and $h/r=0.03$ 1T shown as gray, purple, orange and black lines, respectively. Arrows indicate the force direction: a positive force pushes along $\theta$, while a negative force pushes in the opposite direction. All simulations exhibit similar profiles, showing compression toward the disk midplane above and below it and an expansion region above and below the compression zone. We associate the expansion region with jet expansion and confinement; however, this is not our region of interest. The magnitude of the disk compression force reveals a dichotomy between the thin ($h/r=0.1$ and $h/r=0.03$) and thick ($h/r=0.3$) disks: thicker disks display a stronger compression force above and below the disk midplane. Furthermore, all simulations of thin disks show roughly similar disk compression force magnitudes, even though their disk thicknesses differ by a factor of 3.
@@
We find a stronger disk compression force for thicker disks, because thicker disks restrict the jet expansion, resulting in a stronger backreaction against the disk. When the disk becomes sufficiently thin ($h/r\leq 0.1$), further thinning has minimal impact since the jet occupies most of the available space. The twofold increases in $F_{\mathrm{b,{jet}}}$ compression for thicker vs thinner disks is consistent with the twofold increase in jet efficiency observed in Fig. 6 for thicker vs thinner disks  [^6] The fact that $F_{\mathrm{b,{jet}}}$ is strongest for thick disks may seem at odds with thinner disks being more highly magnetized at the midplane (see Fig. 11). However, the ratio $\frac{1}{2}\langle\delta b^{2}\rangle/(p_{\mathrm{g}}+p_{\mathrm{rad}})$ can be misleading for comparing magnetic field strength across simulations, as thermal (and radiative) pressure also changes with $h/r$. .